For Br2(g), ∆Hfº = 30.9 kJ/mol and ∆Gfº = 3.1 kJ/mol at 298.15 K. Estimate the lowest temperature at which Br2(g) will not spontaneously condense to Br2(l).

A –10. ºC

B –59 ºC

C 59 ºC

D 332 ºC

E 10. ºC
Answer #1

Correct option: C. 59 ºC

---------

∆Hfº = 30.9 kJ/mol

∆Gfº = 3.1 kJ/mol

T = 298.15 K

Now,

∆Gfº = ∆Hfº - T∆Sfº

or, 3.1 = 30.9 - 298.15∆Sfº

or, - 27.8 = - 298.15∆Sfº

or, ∆Sfº = 0.0932 kJ/mol

Now, during phase transition,

∆Gfº = 0

or, ∆Hfº - T∆Sfº = 0

or, 30.9 - T x 0.0932 = 0

or, T x 0.0932 = 30.9

or, T = 332 K

        = (332 - 273) oC

        = 59 oC

Hence, the lowest temperature at which Br2(g) will not spontaneously condense to Br2(l) = 59 oC